# Identify As A Human Does: A Pathfinder of Next-Generation Anti-Cheat Framework for First-Person Shooter Games

### v1

@@ 4.1.1 Temporal Features Construction
    • Offensive props features related with events involving props that aid the attack (e.g., grenade, incendiary, etc.) including the deploy and destroy ticks, the props’ type, the props’ coordinates, etc.
@@
4.1.2 Temporal Data Comparative Visualization
@@
4.2 Structured Features
Structured features are single-value data calculated with well-designed algorithms to represent integrated in-game behavioral information per match. This section describes the components of structured features and visualizes the structured feature differences between normal players and cheaters.
4.2.1 Structured Features Construction
The structured features for Hawk are categorized into four main categories that comprise 28 strong-expressiveness features. For the sake of brevity and coherence, only the introduction of categories is outlined below. A meticulous breakdown can be referred to in Appendix C.
  • Elimination features emphasizes metrics related to in-game eliminations or kill-focused activities. For example, the occluder-penetration index is a weighted index that reflects the extent of the average occluder-penetration eliminations (e.g., eliminate the opponent behind walls or smoke) conducted by the player. In terms of the cheaters exploiting wallhacks, this index is more significant than the normal players’.
Figure 3: Box plots of comparative visualization between honest players and cheaters on top-12 structured features under Mann-Whitney U (Wilcoxon Rank-Sum) test with P-Value ascending (distinction descending) order.
4.2.2 Structured Data Comparative Visualization
Figure 3 offers a box-plot on structured features to analyze between honest players and cheaters comparatively. Each feature on the x-axis (we select top-12 features for illustration purposes) is ordered based on the outcomes of the Mann-Whitney U test, with the leftmost features, such as akpr (introduced in Appendix C), marking pronounced disparities between the two groups. This implies that features towards the left end, showcased by their lower p-values, are more distinguishable, whereas those on the right are less so. The y-axis is normalized values between 0 and 1. The box plots, differentiated by blue for honest players and orange for cheaters, provide insights into the median, interquartile range, and outlier distributions. In which, diamond shapes represent outliers, emphasizing the spread and distribution of data across the features. Despite considering dimensionality reduction techniques, such as PCA (Figure 13), empirical tests show an optimal performance when all 28 features were employed.
4.3 Sense and Performance Features
Sense and performance features are derived from the two subsections above. Appendix D includes detailed classification.
@@ 4.3 Sense and Performance Features
    • Structured performance features contain the remaining structured features that statistically describe the player’s in-game performances, such as time to kill, elimination-related features, etc.
Figure 4: Hawk framework illustration. (a)-RevPov, (b)-RevStats, (c)-ExSPC are corresponding to three observations in Section 3.1, (d)-Mvin is an integration network for combining the determinations from the aforementioned three subsystems.
5 Hawk
To mimic human experts’ identification process, we propose Hawk which consists of four subsystems: Review Point-of-View (RevPov) subsystem, Review Statistics (RevStats) subsystem, Examine Sense-Performance Consistency (ExSPC) subsystem, and Multi-View Integration (Mvin) subsystem. The first three subsystems’ designs are associated with the three observations we discovered regarding the inherent nature of the cheat identification process as elaborated in Section 3.1. The fourth subsystem integrates the results derived from the first three subsystems.
5.1 RevPov
@@
Multi-LSTM Attention Encoders: In Figure 4 (a), we feed the seven sets of temporal features (Section 4.1.1) as input to seven independent Long Short-Term Memory (LSTM) [29] networks to acquire the embeddings. For explanatory purposes, we illustrate the process using the example of the Economy features. $\mathbf{F_{eco}^{n}}$ denotes the input feature vector. A dropout layer [54] is employed from the second to last LSTM layers respectively to mitigate the risk of overfitting. $\vec{h}_{G}^{n}$ denotes the output of the encoder, consists of one set of the seven temporal features mentioned in Section 4.1.1, where $G$ represents the set and $n$ indicates the sequence number among all outputs. For each time step, we introduce the Luong-style attention [41] afterward to help the model better understand the importance of different positions in the input sequence and aggregate different parts of the input weights. In Equation 1, $a_{G}^{i}$ represents the attention weight for the time step $i$. $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ denotes the learned weight matrices of query and key in attention mechanism. $i$, $j$ and $k$ indicate different time steps. Concatenation operation is executed afterward, yielding $[\vec{a}_{G}^{n}||\vec{h}_{G}^{n}]$, where ||||| | denotes concatenation. After two dense layers, $\vec{w}_{G}^{n}$ represents the final tensor for binary classification.
@@
Embedding Generation: An average pooling is conducted to flatten the output sequence. Equation 2 indicates the average operation, where $N$ denotes the number of output sequences. $\vec{V}_{pov}$ is used as embedding for the final determination.
@@
$\vec{V}_{pov}=[\vec{w}_{eco}||\vec{w}_{mov}||\vec{w}_{dmg}||\vec{w}_{elm}||\vec{w}_{aux}||\vec{w}_{off}||\vec{w}_{wf}]$ (3)
Final Determination: In the context of cheat detection, the dataset often exhibits a class imbalance, where the majority of players are honest, inadvertently leading to biased learning towards the dominant class. Therefore, we adopt a multi-subsampling strategy to level the playing field for the minority class, creating balanced representations of both cheating and honest behaviors. Suppose $\mathcal{T}$ is the initial training set. To establish a balanced class distribution, we perform a selective sampling on $\mathcal{T}$, specifically targeting the honest player samples. We perform this operation nine times in accordance with the ratio of honest to dishonest players, generating nine distinct training sets, each denoted as $\mathcal{T}_{k}$, where $k$ indicates the iteration count. Conduct each iteration of the sampling process to select unique honest player samples, thereby adding varied examples to the learning process. Subsequently, we obtain nine separately trained Random Forest models, each corresponding to one of the diversified and balanced training sets $\mathcal{T}_{k}$, we named this strategy as multi-forest. The averaging method in Equation 4 is selected for RevPov final class prediction. Consider a random forest consisting of $N$ decision trees, each producing a prediction denoted as $h_{i}(\vec{V}_{pov}^{k})$, where $i$ indicates the $i$-th tree, $k$ denotes the $k$-th sampled set, and $\vec{V}_{lstm}^{k}$ represents the input feature vector. The prediction of each tree is a probability distribution, indicating the likelihood of the sample $\vec{V}_{pov}^{k}$ belonging to each class. $p_{i}$ represents the probability that the $i$-th tree predicts sample $\vec{V}_{pov}^{k}$ to be in class $j$. In Equation 4, $\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$ denotes the average prediction output of the random forest, which is the average of the prediction probability vectors of all decision trees. The prediction for each $\vec{D}_{k}$ can be obtained by selecting the class with the highest predicted probability shown in Equation 5, where $\left(\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})\right)_{j}$ represents the probability of class $j$ in $\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$. After the multi-forest, we acquire the corresponding determination $\mathbf{D}_{k}$. Lastly, as in Equation 6, we obtain the final classification determination $\mathbf{D}_{pov}$ for RevPov.
@@
5.2 RevStats
Aligned with OBS.2, RevStats subsystem is designed to perform an in-depth analysis of structured features to capture the cheaters. Cheaters often exhibit deviations in their static data compared to normal players. Even when cheaters deliberately attempt to conceal their behavior, the subtle differences tend to accumulate over time and reveal disparities in the statistical data post-game.
Feature Extraction: In Figure 4 (b), $\vec{V}_{28}$ denotes the vector representation of the structured features (Section 4.2.1).
Ensembled Committee: We empirically use seven classifiers [28, 14, 7, 13, 46, 34, 55] in an ensembled way. Each model is independently trained. Upon completion, each model independently adjudicates each test sample. The outputs are binary signifying the presence or absence of cheating.
Final Determination: Due to imbalanced dataset, a similar multi-subsampling strategy as employed in the RevPov subsystem is adopted for the RevStats subsystem to fully leverage the available data without losing valuable information. For each classification model, we generate a distinct model for each subsampled set. Each model provides an individual prediction. A majority voting scheme is utilized within each classification algorithm to aggregate the results of the nine models. The final determination is also achieved through the majority voting mechanism, which collates the determinations from all seven classification algorithms. In Equation 7, where $\mathbf{D}_{stats}$ signifies the binary final determination. Where $w_{j}$ denotes the decision from the $j$th classification algorithm, and $I(\cdot)$ is the indicator function, returning 1 if the condition within the brackets is true, and 0 otherwise. The decision, $i$, which receives the majority of votes from all seven classification algorithms, is accepted as the final determination.
@@ 5.2 RevStats
5.3 ExSPC
Aligned with OBS.3, ExSPC subsystem is designed to inspect a comprehensive investigation of the consistency of a player’s in-game senses and overall performances.
Embedding and Feature Grouping: There are seven embeddings and two groups of vectors as input (Section 4.3) in ExSPC, in which seven embeddings are generated by using the trained models in RevPov and two vectors are the subsets of $\vec{V}_{28}$ in RevStats. Each embedding is padded to the maximum length of that in the whole dataset. $\vec{w}^{k}_{G}$ is the intermediary result from RevPov, where $k$ represents the numbering in the sequence, ${G}$ denotes the type of the temporal features. Each embeddings are represented in two-dimensional arrays. In Figure 4 (c), $P_{1}$ to $P_{k}$ indicates different players, we exemplify with one player $P_{j}$. Therefore, we flatten the embeddings into one-dimensional arrays and use a dense layer with an ELU activation function to shrink the dimensions. After the seven different similar groups of layers, we obtain $\varphi_{G}^{{P}_{j}}$. Subsequently, we divide seven different types of tensors as shown in Appendix D, and concatenate each group of tensors for each player. In which, sense vector of the temporal features is $\vec{V}_{POV}^{S}=[\varphi_{mov}^{{P}_{j}}||\varphi_{eco}^{{P}_{j}}||\varphi_{off}^{{P}_{j}}||\varphi_{aux}^{{P}_{j}}]$ and performance vector of the temporal features is $\vec{V}_{POV}^{P}=[\varphi_{wf}^{{P}_{j}}||\varphi_{elm}^{{P}_{j}}||\varphi_{dmg}^{{P}_{j}}]$. In terms of two vectors, we also divided $\vec{V}_{28}$ into two groups according to Appendix D, namely the sense vector of structured features $\vec{V}_{28}^{S}$ and the performance vector of structured features $\vec{V}_{28}^{P}$.
Dimension Reduction: $\vec{V}_{28}^{S}$ and $\vec{V}_{28}^{P}$ are normalized at first. Afterward, each group of the four vectors goes through different groups of dense layers to reduce the dimensions. Next, we concatenate four vectors and yield two vectors eventually representing sense $\vec{V}_{SENSE}=[\vec{\xi}_{SENSE}^{STATS}||\vec{\xi}_{SENSE}^{POV}]$ and performance $\vec{V}_{PREF}=[\vec{\xi}_{PREF}^{STATS}||\vec{\xi}_{PREF}^{POV}]$ respectively.
Final Determination: When acquiring $\vec{V}_{SENSE}$ and $\vec{V}_{PERF}$, we further reduce its dimension and deepen the network for better learning performances. Then, we concatenate two vectors and go through one final dense layer for reducing the dimension and another dense layer with sigmoid activation function for binary classification. Eventually, the final classification determination $\mathbf{D}_{spc}$ for ExSPC subsystem is obtained. The class weight is set as one to nine like previous subsystems. Both RevPov and ExSPC use Binary Cross-Entropy as loss function shown in Equation 8, where $N$ represents the number of samples, $y_{i}$ is the true label of the $i$-th sample, $\hat{y}_{i}$ is the predicted value for the $i$-th sample.
@@ 5.3 ExSPC
5.4 Mvin
In Figure 4 (d), Mvin is to harness the collective determinations from RevPov, RevStats, and ExSPC. By integrating the predictions from these subsystems, Mvin aims to deliver a final verdict regarding a player’s cheating disposition.
@@ 5.4 Mvin
[…]
6 Experiments
We conduct experiments to answer six research questions (RQ.1--RQ.6):
  • RQ.1: What are the performances of Hawk and each subsystem? (Section 6.4)
  • RQ.2: Compared to the current official inspection, to what extent Hawk has improved in terms of effectiveness and efficiency? (Section 6.5)
  • RQ.3: What are the performances of Hawk under different settings of the Task-Specified Threshold Optimizer? (Section 6.6)
  • RQ.4: What is the robustness of Hawk and its subsystems against cheat evolution? (Section 6.7)
  • RQ.5: Is there any suspicious activity that still exists but has never been detected by officials? (Section 6.8)
  • RQ.6: What are the overheads of Hawk’s subsystems? (Section 6.9)
6.1 Incomparability to Prior Works
First, most prior works [69, 24, 70, 2, 39, 71] are not open-sourced. Second, data collection method is distinct and incompatible. Hawk utilizes the replay file post-game, which differs fundamentally from prior studies that modified the game engine or applied mods during gameplay. Third, the installation on commercial clients or servers is infeasible. If the installation of the prior works is implemented before the commencement of data collection, the data can then be gathered and evaluated. However, the demos we collected come directly from the game platform, which prohibits the installation of any unverified programs on clients or servers due to legal constraints. In addition to the aforementioned reasons with commonality, we also further state below the specific reasons for the detailed incomparability between Hawk and the SOTA designs. There are three aspects of incomparability with BotScreen [9]. First, CS:GO has gone through several major patch updates after BotScreen was released. The authors of BotScreen confirmed through email that BotScreen, as well as Osiris [37] (the identical cheat used in BotScreen for feature extraction and evaluation), does not apply to the in-use version of games. Second, the deprecation of Intel SGX [5, 31] prevents us from conducting comparative experiments. Third, it can only detect pure aimbot, making the comparison less significant. BlackMirror [49] and Invisibility Cloak [58] are prevention designs only for counteracting wallhacks and computer-vision-based aimbots respectively. Both studies are only functional before the cheats happen. Whereas Hawk and other prior works are detection systems that functional post-cheating.
Table 1: Dataset separation and description.

### v2

@@ IV-B1 Temporal Features Construction
    • Offensive props features related with events involving props that aid the attack (e.g., grenade, incendiary, etc.) including the deploy and destroy ticks, the props’ type and coordinates, etc.
@@
IV-B2 Temporal Data Comparative Visualization
@@
IV-C Sense and Performance Features
Sense and performance features are derived from the previous subsections. Supplemental Material E includes detailed classification.
@@ IV-C Sense and Performance Features
    • Structured performance features contain the remaining structured features that statistically describe the player’s in-game performances, e.g., elimination-related features.
Figure 4: Hawk framework illustration. (a)-RevPov, (b)-RevStats, (c)-ExSPC are corresponding to three observations in Section III-A, (d)-Mvin is an integration network for combining the determinations from the aforementioned three subsystems.
V Hawk
To mimic human experts’ identification process, we propose Hawk which consists of four subsystems: Review Point-of-View (RevPov), Review Statistics (RevStats), Examine Sense-Performance Consistency (ExSPC), and Multi-View Integration (Mvin). The first three subsystems are associated with the three observations we discovered regarding the inherent nature of the cheat identification process in Section III-A. Mvin integrates the results derived from the above three subsystems, just as humans integrate the three aspects collectively.
V-A RevPov
@@
Multi-LSTM Attention Encoders: In Figure 4a, we feed the seven sets of temporal features (Section IV-B1) as input to seven independent Long Short-Term Memory (LSTM) [42] networks to acquire the embeddings. For explanatory purposes, we illustrate the process using the example of the Economy features. $\mathbf{F_{eco}^{n}}$ denotes the input feature vector. Dropout layers [43] are employed to mitigate the risk of overfitting. $\vec{h}_{G}^{n}$ denotes the output of the encoder, consists of one set of the seven temporal features mentioned in Section IV-B1, where $G$ represents the set and $n$ indicates the sequence number among all outputs. For each time step, we introduce the Luong-style attention [44] afterward to help the model better understand the importance of different positions in the input sequence and aggregate different parts of the input weights, overcoming the challenge of identifying varying patterns in temporal data for a large number of samples. In Equation 1, $a_{G}^{i}$ represents the attention weight for the time step $i$. $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ denotes the learned weight matrices of query and key in attention mechanism. $i$, $j$ and $k$ indicate different time steps. Concatenation operation is executed afterward, yielding $\vec{a}_{G}^{n}||\vec{h}_{G}^{n}]$, where $||$ denotes concatenation. After two dense layers, $\vec{w}_{G}^{n}$ represents the final tensor for binary classification.
@@
Embedding Generation: An average pooling in Equation 2 is conducted to flatten the output sequence, where $N$ denotes the number of output sequences. $\vec{V}_{pov}$ is the final determination embedding.
@@
[…]
Final Determination: We find that only depending on the fully connected layers with different activation functions can not effectively classify. Therefore, we conduct the following strategies to further improve RevPov. In the context of cheat detection, the dataset often exhibits a class imbalance, where the majority of players are honest, inadvertently leading to biased learning towards the dominant class. Therefore, we create a multi-subsampling strategy to level the playing field for the minority class, creating balanced representations of both cheating and honest behaviors without dropping any samples. We perform a selective sampling to divide the honest players in the initial training set $\mathcal{T}$ into nine sets in accordance with the ratio of normal to cheating players, generating nine training sets containing the same set of cheaters and distinct normal players, each denoted as $\mathcal{T}_{k}$, where $k$ indicates the iteration count. Subsequently, after adding varied examples to the learning process, we obtain nine independent Random Forest models, each corresponding to one of the diversified and balanced training sets $\mathcal{T}_{k}$, we named this strategy as multi-forest. The averaging method in Equation 4 is selected for RevPov final class prediction, where random forest consisting of $N$ decision trees, each producing a prediction $h_{i}(\vec{V}_{pov}^{k})$, where $i$ indicates the $i$-th tree, $k$ denotes the $k$-th sampled set, and $\vec{V}_{pov}^{k}$ represents the input feature vector. $p_{i}$ represents the probability that the $i$-th tree predicts sample $\vec{V}_{pov}^{k}$ to be in class $j$. $\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$ denotes the average prediction output of the random forest, which is the average of the prediction probability vectors of all decision trees. The prediction for each $\vec{D}_{k}$ is obtained by selecting the class with the highest predicted probability shown in Equation 5, where $(\cdot)_{j}$ represents the probability of class $j$ in $\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$. After the multi-forest, we acquire the corresponding determination $\mathbf{D}_{k}$. Lastly, as in Equation 6, we obtain the final classification determination $\mathbf{D}_{pov}$ for RevPov.
@@
V-B RevStats
Aligned with OBS.2, RevStats subsystem is designed to perform an in-depth analysis of structured features to capture the cheaters.
Feature Extraction: In Figure 4b, $\vec{V}_{28}$ denotes the vector representation of the structured features (Section IV-A1).
Ensembled Committee: During preliminary testing, our selected features showed promising results across most classifiers, demonstrating their effectiveness. However, despite similar performance in terms of metrics alone, the testing revealed that different classifiers produced varying classifications on some individual samples. This suggests that relying on a single classifier, as done in prior works, can lead to unreliable outcomes. However, the majority of classifiers correctly classified the samples, leading to the development of the Ensembled Committee. Among hundreds of classifiers, we identify seven superior classifiers [45, 46, 47, 48, 49, 50, 51] after the preliminary testing and use them in an ensembled way.
Final Determination: Due to imbalanced dataset, a similar multi-subsampling strategy as employed in the RevPov subsystem is adopted for the RevStats subsystem to fully leverage the available data without losing valuable information. For each classification model, we generate a distinct model for each subsampled set. Each model is independently trained and provides an individual prediction. A majority voting scheme is utilized within each classification algorithm to aggregate the results of the seven models. The final determination is also achieved through majority voting, which collates the determinations from all nine subsampled sets. In Equation 7, where $\mathbf{D}_{stats}$ signifies the binary final determination. Where $w_{j}$ denotes the decision from the $j$th classification algorithm, and $I(\cdot)$ is the indicator function, returning 1 if the condition within the brackets is true, and 0 otherwise. The decision, $i$, which receives the majority of votes from all seven classification algorithms, is accepted as the final determination.
@@ V-B RevStats
V-C ExSPC
Aligned with OBS.3, ExSPC is designed from scratch to inspect the consistency of a player’s in-game senses and overall performances.
Embedding and Feature Grouping: ExSPC takes seven embeddings and two groups of vectors as input (see Section IV-C). The seven embeddings are generated using trained models from RevPov, while the two vectors are subsets of $\vec{V}_{28}$ from RevStats. Each embedding is padded to the maximum length observed in the dataset. The intermediary result from RevPov, denoted as $\vec{w}^{k}_{G}$, represents temporal features, where $k$ is the sequence number and G{G} denotes the type of the temporal features. All embeddings are structured as two-dimensional arrays. As illustrated in Figure 4c, where $P_{1}$ to $P_{k}$ represent different players, we exemplify with one player $P_{j}$. The embeddings are first flattened and dimensionally reduced. After processing through seven groups of similar layers, we obtain $\varphi_{G}^{{P}_{j}}$. These tensors are then divided into seven types and concatenated for each player. In which, $\vec{V}_{POV}^{S}$ denotes the temporal sense features vector and $\vec{V}_{POV}^{P}$ denotes the temporal performance features vector. Similarly, the structured features vector $\vec{V}_{28}$ is divided into two groups: the sense vector $\vec{V}_{28}^{S}$ and the performance vector $\vec{V}_{28}^{P}$, as detailed in Supplemental Material E.
Dimension Reduction: $\vec{V}_{28}^{S}$ and $\vec{V}_{28}^{P}$ are normalized. Afterward, each group of the four vectors goes through different groups of dense layers to reduce the dimensions. Next, concatenate four vectors and yield two vectors eventually representing sense $\vec{V}_{SENSE}$ and performance $\vec{V}_{PREF}$ respectively.
Final Determination: When acquiring $\vec{V}_{SENSE}$ and $\vec{V}_{PERF}$, we further reduce their dimensions and deepen the network for better learning performances. Then, we concatenate two vectors and go through one final dense layer for reducing the dimension and another dense layer with sigmoid activation function for binary classification. Eventually, the final classification determination $\mathbf{D}_{spc}$ for ExSPC subsystem is obtained. The class weight is set as one to nine like previous subsystems. Both RevPov and ExSPC use Binary Cross-Entropy as loss function shown in Equation 8, where $N$ represents the number of samples, $y_{i}$ is the true label of the $i$-th sample, $\hat{y}_{i}$ is the predicted value for the $i$-th sample.
@@ V-C ExSPC
V-D Mvin
In Figure 4d, Mvin is to dynamically integrate determinations from RevPov, RevStats, and ExSPC.
@@ V-D Mvin
Model Optimization: Mvin adopts a weight matrix for each subsystem’s determination, where weights are dynamically assigned based on the subsystem’s importance and reliability. The weighted output, $\mathbb{W}$, is calculated as Equation 9. Where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ represent the weights associated with the outputs of the subsystems RevPov, RevStats, and ExSPC, respectively.
$\mathbb{W}=\lambda_{1}\cdot\mathbf{D}_{pov}+\lambda_{2}\cdot\mathbf{D}_{stats}+\lambda_{3}\cdot\mathbf{D}_{spc}$ (9)
Final Determination: $\mathbb{W}$ undergoes a learnable thresholding via the Task-Specified Threshold Optimizer (TSTO), which provides a binary determination based on a dynamic threshold $\varepsilon$. The threshold optimizer can be dynamically set based on anti-cheat requirements. For instance, in Recall mode, the optimizer would aim to identify every potential cheater, accepting the risk of higher false positives. This threshold $\varepsilon$ therefore acts as a decision boundary, converting the continuous weighted value $\mathbb{W}$ into a discrete binary outcome $\mathbf{D}_{HAWK}$. TSTO offers Hawk a degree of selectivity, enabling flexible tuning and offering promising performance without requiring retraining any prior subsystems, overcoming the challenge of lack of system adaptability. By harnessing the strengths of each subsystem and compensating for their limitations, Mvin is crucial to integrate determinations from different subsystems.
VI Experiments
We conduct experiments to answer seven research questions (RQ.1-RQ.7):
  • RQ.1: What are the performances of Hawk and each subsystem? (Section VI-C)
  • RQ.2: To what extent has Hawk improved effectiveness and efficiency compared to current official inspections? (Section VI-D)
  • RQ.3: What are the performances of Hawk under different settings of the Task-Specified Threshold Optimizer? (Section VI-F)
  • RQ.4: What is the robustness of Hawk and its subsystems against cheat evolution? (Section VI-G)
  • RQ.5: Is there any suspicious activity that still exists but has never been detected by officials? (Section VI-H)
  • RQ.6: What is the overhead of Hawk? (Section VI-I)
  • RQ.7: Is Hawk performing better than prior works? (Section VI-E)
Figure 5: Experiments’ process and labeling details.
TABLE II: Dataset separation and description.